# Entanglement Hamiltonian and effective temperature of non-Hermitian quantum spin ladders

### v1

@@ Abstract
Quantum entanglement plays a crucial role not only in understanding Hermitian many-body systems but also in offering valuable insights into non-Hermitian quantum systems. In this paper, we analytically investigate the entanglement Hamiltonian and entanglement energy spectrum of a non-Hermitian spin ladder using perturbation theory in the biorthogonal basis. Specifically, we examine the entanglement properties between coupled non-Hermitian quantum spin chains. In the strong coupling limit ($J_{\mathrm{rung}}\gg 1$), first-order perturbation theory reveals that the entanglement Hamiltonian closely resembles the single-chain Hamiltonian with renormalized coupling strengths, allowing for the definition of an ad hoc temperature. Our findings provide new insights into quantum entanglement in non-Hermitian systems and offer a foundation for developing novel algorithms, such as applying finite-temperature Density Matrix Renormalization Group (DMRG) to non-Hermitian quantum systems.
I Introduction
@@
Interesting phenomena related to the entanglement Hamiltonian $H_{E}$ can also arise when the subsystem B is considered as an ancilla system copied from the subsystem A, and introducing strong enough interaction $H_{AB}$ to create a nearly maximally entangled state between A and B. [10, 11, 12] For example, in an antiferromagnetic spin ladder where the rung coupling is much stronger than the leg coupling, the ground-state forms multiple rung-singlets, resulting in a nearly maximally entangled state between the two legs. [13, 14, 15, 16, 17, 18, 19, 20, 21] In contrast to the Li-Haldane conjecture in topological systems, the entire entanglement spectrum has some similarity to the energy spectrum of subsystem A. Remarkably, under carefully selected parameters, the entanglement Hamiltonian $H_{E}\approx\beta H_{A}$ can be proportional to the Hamiltonian of subsystem A by a constant $\beta$. In other words, the finite temperature properties of an isolated system A can be approximated by the reduced density matrix $\rho_{A}$ obtained from the ground-state of an enlarged system at zero temperature,
@@
where $\beta$ is the inverse temperature as a function of system parameters. This development allows the finite-temperature Density Matrix Renormalization Group (DMRG) based on matrix product states to evolve into a robust numerical method for finite-temperature strongly correlated systems. [10]
@@
In this paper, we show that Eq. (1) remains valid for non-Hermitian Hamiltonians. This implies that the ancilla trick used in finite-temperature algorithms [10, 11, 12] is also expected to be effective for non-Hermitian quantum systems. In the next section, we briefly introduce non-Hermitian quantum mechanics and describe the non-Hermitian Hamiltonian for the spin-1/2 ladder.
II non-Hermitian Hamiltonian
@@
Non-Hermitian systems have become an important multidisciplinary field of study, [22, 23, 24] spanning photonics, [25] condensed matter physics, [26, 27] and quantum information science. [28, 29, 30, 31, 32, 33, 34, 35, 36] Unlike traditional quantum systems, which are governed by Hermitian Hamiltonians that ensure real eigenvalues and physically observable energy levels, non-Hermitian systems are described by Hamiltonians that do not necessarily satisfy this condition. This leads to complex eigenvalues and unique physical phenomena, e.g. exceptional points [37, 38, 39, 40, 41, 42] and non-Hermitian skin effect. [43, 44, 45] An exceptional point (EP) in non-Hermitian systems occurs when two or more eigenvalues and their corresponding eigenvectors coalesce, rendering the Hamiltonian non-diagonalizable. This leads to unique phenomena, such as enhanced sensitivity to external changes [25] and a negative divergence in real part of fidelity susceptibility [39, 40, 41] or quantum metric, [46, 47] which measures how ground state changes under perturbations. The non-Hermitian skin effect arises in systems with nonreciprocal coupling. [43] This nonreciprocity means that particles or excitations have different probabilities of hopping forward versus backward. As a result, even in the absence of an external field or disorder, the system can exhibit an accumulation of states at one end under open boundary condition, leading to the Hermitian bulk – non-Hermitian boundary correspondence. [48]
@@
Mathematically, nonreciprocal coupling is often introduced into a system’s Hamiltonian through asymmetric hopping terms. [26, 27] For instance, in a one-dimensional lattice model, the hopping amplitude from site $j$ to site $j+1$ might differ from the amplitude from site $j+1$ to site $j$. This asymmetry results in a complex band structure with eigenvalues that can form loops in the complex plane, leading to the skin effect. [48] In this paper, we study the following non-Hermitian Hamiltonian $H=H_{A}+H_{B}+H_{AB}$ for spin-1/2 ladder with the nonreciprocal coupling.
@@
where $N$ denotes the number of rungs, and $\Phi$ and $\Psi$ are real parameters that control the nonreciprocal coupling between the legs and within each leg, respectively. $J_{\mathrm{rung}}$ and $J_{\mathrm{leg}}$ represent the coupling strengths between the legs and within the legs. Lastly, $\Delta$ denotes the XXZ anisotropy strength. Periodic boundary conditions are assumed. In the Hermitian limit, $\Phi=\Psi=0$, the ground-state phase diagram has been studied in the literature. [49] We will mainly focus on the entanglement Hamiltonian in the rung-singlet phase. The schematic representation of the non-Hermitian spin-1/2 ladder is shown in Fig. 1.
@@
Due to the non-Hermitian nature of the system, where $H^{\dagger}\neq H$, the time evolution of the wavefunctions is driven by both $H$ and $H^{\dagger}$ simultaneously. This results in the following time evolution equations: $\frac{\partial}{\partial t}|\varphi^{R}(t)\rangle{=}{-}iH|\varphi^{R}(t)\rangle$, and $\frac{\partial}{\partial t}|\varphi^{L}(t)\rangle{=}{-}iH^{\dagger}|\varphi^{L}(t)\rangle$. Note that $\hslash\equiv 1$ is set. Consequently, in analogy to standard linear algebra, the eigenvectors of a non-Hermitian Hamiltonian are generalized into biorthogonal left and right eigenvectors, satisfying the following eigenvalue equations: $H^{\dagger}|\psi_{n}^{L}\rangle=E_{n}^{*}|\psi_{n}^{L}\rangle$ and $H|\psi_{n}^{R}\rangle=E_{n}|\psi_{n}^{R}\rangle$, with the biorthonormal condition: $\langle\psi_{n}^{L}|\psi_{m}^{R}\rangle{=}\delta_{nm}$[24] In non-Hermitian quantum mechanics, observables are defined through the expectation value $\langle\psi^{L}|O|\psi^{R}\rangle$, involving both left and right eigenvectors. This biorthogonal framework naturally extends to the definition of the reduced density matrix in a bipartite system. For a system divided into subsystems A and B, the reduced density matrix (RDM) of the ground-state for subsystem A is defined as
@@
III Entanglement Hamiltonian
III.1 Perturbation Theory
@@
$=\frac{J_{\mathrm{leg}}}{4J_{\mathrm{rung}}}\sum_{j=1}^{N}\Big{[}\frac{2e^{-\Phi}e^{-\Psi}}{1+\Delta}...\langle t_{j}^{+}|\langle t_{j+1}^{-}|...$
$+\frac{2e^{-\Phi}e^{\Psi}}{1+\Delta}...\langle t_{j}^{-}|\langle t_{j+1}^{+}|...-\Delta...\langle t_{j}^{0L}|\langle t_{j+1}^{0L}|...\Big{]},$ (10a)
@@
$=\frac{J_{\mathrm{leg}}}{4J_{\mathrm{rung}}}\sum_{j=1}^{N}\Big{[}\frac{2e^{\Phi}e^{\Psi}}{1+\Delta}...|t_{j}^{+}\rangle|t_{j+1}^{-}\rangle...$
$+\frac{2e^{\Phi}e^{-\Psi}}{1+\Delta}...|t_{j}^{-}\rangle|t_{j+1}^{+}\rangle...-\Delta...|t_{j}^{0R}\rangle|t_{j+1}^{0R}\rangle...\Big{]}.$ (10b)
@@
$\rho_{A}$ $\approx\rho_{A}^{(0)}+\rho_{A}^{(1)}$
$=\mathrm{Tr}_{B}\left(|\psi_{0}^{R(0)}\rangle\langle\psi_{0}^{L(0)}|+|\psi_{0}^{R(1)}\rangle\langle\psi_{0}^{L(0)}|+|\psi_{0}^{R(0)}\rangle\langle\psi_{0}^{L(1)}|\right)$
$=\frac{1}{2^{N}}\Big{(}1-\frac{4J_{\mathrm{leg}}}{J_{\mathrm{rung}}(1+\Delta)}\sum_{j=1}^{N}\left[\frac{1}{2}\Big{(}e^{\Psi}S_{j,A}^{+}S_{j+1,A}^{-}\right.$
$\left.+e^{-\Psi}S_{j,A}^{-}S_{j+1,A}^{+}\Big{)}+\frac{1}{2}\left(\Delta+\Delta^{2}\right)S_{j,A}^{z}S_{j+1,A}^{z}\right]\Big{)}$ (13)
@@
$\rho_{A}$ $\approx\frac{1}{Z}\exp[-\beta\tilde{H}_{A}]$
$=\frac{1}{Z}\left(1-\beta\tilde{H}_{A}+\frac{1}{2!}\beta^{2}\tilde{H}_{A}^{2}-\frac{1}{3!}\beta^{3}\tilde{H}_{A}^{3}+\cdots\right)$ (14)
Compare Eq. (III.1) with Eq. (III.1), we obtain the ad hoc inverse temperature
@@
$\tilde{H}_{A}=$ $J_{\mathrm{leg}}\sum_{j=1}^{N}\Big{[}\frac{1}{2}\left(e^{\Psi}S_{j,A}^{+}S_{j+1,A}^{-}+e^{-\Psi}S_{j,A}^{-}S_{j+1,A}^{+}\right)$
$+\tilde{\Delta}S_{j,A}^{z}S_{j+1,A}^{z}\Big{]},$ (16)
@@
III.2 Discussion
We make some remarks regarding the derivation: For the spin ladder Eq. (2) considered in this paper, the renormalized anisotropy parameter remains unchanged, i.e., $\tilde{\Delta}=\Delta$, when $\Delta=1$ or 00. In these specific cases, the entanglement Hamiltonian is exactly equal to the subsystem Hamiltonian, $\tilde{H}_{A}=H_{A}$. When $\Psi=\Phi=0$, our results are consistent with those from earlier studies [16], reproducing the known Hermitian case. Even when non-Hermitian couplings are introduced with non-zero $\Psi$ and $\Phi$, the overall behavior remains remarkably similar, confirming that the methods and conclusions from the Hermitian regime can be successfully extended to non-Hermitian systems without major deviations. In the general non-Hermitian case, if one wishes to ensure that $\tilde{H}_{A}=H_{A}$, a simple approach is to choose the inter-subsystem Hamiltonian $H_{AB}$ as a Hermitian and isotropic Heisenberg interaction.
@@ IV Conclusion
In this paper, we explore the entanglement Hamiltonian of a non-Hermitian spin ladder system using perturbation theory, providing an analytical approach to this difficult problem. Remarkably, we find that the entanglement Hamiltonian in the non-Hermitian case can be approximated by the Hamiltonian of subsystem A, indicating that the thermal density matrix of an isolated non-Hermitian system A in equilibrium can be derived by the partial trace of an enlarged system. This suggests that the ancilla trick applied for developing finite-temperature Density Matrix Renormalization Group (DMRG) method [10], can be extended to non-Hermitian many-body systems as well. Our work offers new insights into the study of quantum entanglement in non-Hermitian systems, potentially facilitating the development of advanced numerical algorithms for investigating their finite-temperature behavior.
Although non-Hermitian systems exhibit many phenomena absent in Hermitian systems, such as exceptional points (EP) and the non-Hermitian skin effect, many features of Hermitian systems persist in non-Hermitian counterparts when appropriately generalized. For instance, the entanglement entropy in critical systems still follows a logarithmic scaling [28, 29], fidelity susceptibility diverges near phase transitions or EPs [39, 40], and machine learning methods can be transferred from Hermitian to non-Hermitian systems [53]. In our work on entanglement Hamiltonians, we extend the Hermitian case to non-Hermitian systems and find that, for nearly maximally entangled states, the results remain consistent with the Hermitian case.
Acknowledgements.
YCT is grateful to Chia-Yi Ju, Po-Yao Chang and Gunnar Möller for many invaluable discussions. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources. YCT is grateful to the supports from National Science and Technology Council (NSTC) of Taiwan under grant No. 113-2112-M-A49-015-MY3.

### v2

@@ Abstract
Quantum entanglement plays a crucial role not only in understanding Hermitian many-body systems but also in offering valuable insights into non-Hermitian quantum systems. In this paper, we analytically investigate the entanglement Hamiltonian and entanglement energy spectrum of a non-Hermitian spin ladder using perturbation theory in the biorthogonal basis. Specifically, we examine the entanglement properties between coupled non-Hermitian quantum spin chains. In the strong coupling limit ($J_{\mathrm{rung}}\gg 1$), first-order perturbation theory reveals that the entanglement Hamiltonian closely resembles the single-chain Hamiltonian with renormalized coupling strengths, allowing for the definition of an ad hoc temperature. Our findings provide new insights into quantum entanglement in non-Hermitian systems and offer a foundation for developing novel approaches for studying finite temperature properties in non-Hermitian quantum many-body systems.
Copyright attribution to authors.
This work is a submission to SciPost Physics Core.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
Interesting phenomena related to the entanglement Hamiltonian $H_{E}$ can also arise when subsystem B is considered a copy of subsystem A, and a strong enough interaction $H_{AB}$ is introduced to create a nearly maximally entangled ground-state between A and B. For example, in an antiferromagnetic spin ladder where the rung coupling is much stronger than the leg coupling, the ground-state forms multiple rung-singlets, resulting in a nearly maximally entangled state between the two legs[10, 11, 12, 13, 14, 15, 16, 17, 18]. In contrast to the Li-Haldane conjecture in topological systems, the entire entanglement spectrum has some similarity to the energy spectrum of subsystem A. Remarkably, under carefully selected parameters, the entanglement Hamiltonian $H_{E}\approx\beta H_{A}$ can be proportional to the Hamiltonian of subsystem A by a constant $\beta$. In other words, the finite temperature properties of an isolated system A can be approximated by the reduced density matrix $\rho_{A}$ obtained from the ground-state of an enlarged system at zero temperature,
@@
where $\beta$ is the inverse temperature as a function of system parameters.
It is worth noting that the finite-temperature Density Matrix Renormalization Group (DMRG) is a well-established numerical method for obtaining the finite-temperature properties of a system [19]. This method relies on the ancilla trick [20], which involves pairing the original system with an auxiliary system, known as the ancilla, to create an enlarged system. In this framework, the initial state is intentionally selected as a maximally entangled pure state shared between the original system and the ancilla. This state represents the system at infinite temperature. By performing imaginary time evolution on this initial state, the system can be effectively cooled to finite temperatures. This process enables the calculation of thermal expectation values.
In contrast, Eq. (1) provides an alternative approach that avoids imaginary time evolution. By examining the ground state of an enlarged system, it is possible to directly approximate the original system’s finite-temperature properties, offering a simpler method for studying finite-temperature behavior from ground-state entanglement alone. This approach could also inspire experimental designs for simulating finite-temperature properties in real systems.
@@
In this paper, we show that Eq. (1) remains valid for non-Hermitian Hamiltonians. In the next section, we briefly introduce non-Hermitian quantum mechanics and describe the non-Hermitian Hamiltonian for the spin-1/2 ladder.
2 non-Hermitian Hamiltonian
@@
Non-Hermitian systems have become an important multidisciplinary field of study [21, 22, 23], spanning photonics [24], condensed matter physics [25, 26, 27, 28, 29], and quantum information science. [30, 31, 32, 33, 34, 35, 36, 37, 38] A novel approach to studying non-Hermitian systems on a quantum computer can be achieved by applying post-selection to qubits, enabling non-reciprocal and non-unitary evolution through controlled qubit operations [39]. Unlike traditional quantum systems, which are governed by Hermitian Hamiltonians that ensure real eigenvalues and physically observable energy levels, non-Hermitian systems are described by Hamiltonians that do not necessarily satisfy this condition. This leads to complex eigenvalues and unique physical phenomena, e.g. exceptional points [40, 41, 42, 43, 44, 45, 46] and non-Hermitian skin effect. [47, 48, 49, 50, 51, 52, 53, 54] An exceptional point (EP) in non-Hermitian systems occurs when two or more eigenvalues and their corresponding eigenvectors coalesce, rendering the Hamiltonian non-diagonalizable. This leads to unique phenomena, such as enhanced sensitivity to external changes [24] and a negative divergence in real part of fidelity susceptibility [42, 43, 44] or quantum metric, [55, 56, 57] which measures how ground state changes under perturbations. The non-Hermitian skin effect refers to the phenomenon where a macroscopic fraction of the system’s eigenstates accumulate at one edge under open boundary conditions, even in the absence of external fields or disorder [47]. This effect is also commonly observed in systems with nonreciprocal coupling, where particles or excitations have different probabilities of hopping forward versus backward [48, 49].
@@
Mathematically, nonreciprocal coupling is often introduced into a system’s Hamiltonian through asymmetric hopping terms. [25, 26, 27] For instance, in a one-dimensional lattice model, the hopping amplitude from site $j$ to site $j+1$ might differ from the amplitude from site $j+1$ to site $j$. This asymmetry results in a complex band structure with eigenvalues that can form loops in the complex plane, leading to the skin effect [53]. In this paper, we study the following non-Hermitian Hamiltonian $H=H_{A}+H_{B}+H_{AB}$ for spin-1/2 ladder with the nonreciprocal coupling.
@@
where $N$ denotes the number of rungs, and $\Phi$ and $\Psi$ are real parameters that control the nonreciprocal coupling between the legs and within each leg, respectively. $J_{\mathrm{rung}}$ and $J_{\mathrm{leg}}$ represent the coupling strengths between the legs and within the legs. Lastly, $\Delta$ denotes the XXZ anisotropy strength. Periodic boundary conditions are assumed. In the Hermitian limit, $\Phi=\Psi=0$, the ground-state phase diagram has been studied in the literature [58]. We will mainly focus on the entanglement Hamiltonian in the rung-singlet phase. The schematic representation of the non-Hermitian spin-1/2 ladder is shown in Fig. 1.
@@
[…]
@@
3 Entanglement Hamiltonian
3.1 Perturbation Theory
@@
$=\frac{J_{\mathrm{leg}}}{4J_{\mathrm{rung}}}\sum_{j=1}^{N}\left[\frac{2e^{-\Phi}e^{-\Psi}}{1+\Delta}...\langle t_{j}^{+}|\langle t_{j+1}^{-}|...+\frac{2e^{-\Phi}e^{\Psi}}{1+\Delta}...\langle t_{j}^{-}|\langle t_{j+1}^{+}|...-\Delta...\langle t_{j}^{0L}|\langle t_{j+1}^{0L}|...\right],$ (10a)
@@
$=\frac{J_{\mathrm{leg}}}{4J_{\mathrm{rung}}}\sum_{j=1}^{N}\left[\frac{2e^{\Phi}e^{\Psi}}{1+\Delta}...|t_{j}^{+}\rangle|t_{j+1}^{-}\rangle...+\frac{2e^{\Phi}e^{-\Psi}}{1+\Delta}...|t_{j}^{-}\rangle|t_{j+1}^{+}\rangle...-\Delta...|t_{j}^{0R}\rangle|t_{j+1}^{0R}\rangle...\right].$ (10b)
@@
$\rho_{A}$ $\approx\rho_{A}^{(0)}+\rho_{A}^{(1)}=\mathrm{Tr}_{B}\left(|\psi_{0}^{R(0)}\rangle\langle\psi_{0}^{L(0)}|+|\psi_{0}^{R(1)}\rangle\langle\psi_{0}^{L(0)}|+|\psi_{0}^{R(0)}\rangle\langle\psi_{0}^{L(1)}|\right)$
$=\frac{1}{2^{N}}\left(1-\frac{4J_{\mathrm{leg}}}{J_{\mathrm{rung}}(1+\Delta)}\sum_{j=1}^{N}\left[\frac{1}{2}\left(e^{\Psi}S_{j,A}^{+}S_{j+1,A}^{-}+e^{-\Psi}S_{j,A}^{-}S_{j+1,A}^{+}\right)+\frac{1}{2}\left(\Delta+\Delta^{2}\right)S_{j,A}^{z}S_{j+1,A}^{z}\right]\right)$ (13)
@@
$\rho_{A}\approx\frac{1}{Z}\exp[-\beta\tilde{H}_{A}]=\frac{1}{Z}\left(1-\beta\tilde{H}_{A}+\frac{1}{2!}\beta^{2}\tilde{H}_{A}^{2}-\frac{1}{3!}\beta^{3}\tilde{H}_{A}^{3}+\cdots\right)$ (14)
Compare Eq. (14) with Eq. (3.1), we obtain the ad hoc inverse temperature
@@
$\tilde{H}_{A}=J_{\mathrm{leg}}\sum_{j=1}^{N}\left[\frac{1}{2}\left(e^{\Psi}S_{j,A}^{+}S_{j+1,A}^{-}+e^{-\Psi}S_{j,A}^{-}S_{j+1,A}^{+}\right)+\tilde{\Delta}S_{j,A}^{z}S_{j+1,A}^{z}\right],$ (16)
@@
3.2 Discussion
We make some remarks regarding the derivation: For the spin ladder Eq. (2) considered in this paper, the renormalized anisotropy parameter remains unchanged, i.e., $\tilde{\Delta}=\Delta$, when $\Delta=1$ or $0$. In these specific cases, the entanglement Hamiltonian is exactly equal to the subsystem Hamiltonian, $\tilde{H}_{A}=H_{A}$. When $\Psi=\Phi=0$, our results are consistent with those from earlier studies [13], reproducing the known Hermitian case. Even when non-Hermitian couplings are introduced with non-zero $\Psi$ and $\Phi$, the overall behavior remains remarkably similar, confirming that the methods and conclusions from the Hermitian regime can be successfully extended to non-Hermitian systems without major deviations. In the general non-Hermitian case, if one wishes to ensure that $\tilde{H}_{A}=H_{A}$, a simple approach is to choose the inter-subsystem Hamiltonian $H_{AB}$ as a Hermitian and isotropic Heisenberg interaction.
@@ 4 Conclusion
In this paper, we explore the entanglement Hamiltonian of a non-Hermitian spin ladder system using perturbation theory, providing an analytical approach to this difficult problem. Remarkably, we find that the entanglement Hamiltonian in the non-Hermitian case can be approximated by the Hamiltonian of subsystem A, indicating that the thermal density matrix of an isolated non-Hermitian system A can be derived by the partial trace of an enlarged system. Our work offers new insights into the study of quantum entanglement in non-Hermitian systems, potentially facilitating the development of advanced approaches for investigating their finite-temperature behavior.
Although non-Hermitian systems exhibit many phenomena absent in Hermitian systems, such as exceptional points (EP) and the non-Hermitian skin effect, many features of Hermitian systems persist in non-Hermitian counterparts when appropriately generalized. For instance, the entanglement entropy in critical systems still follows a logarithmic scaling [30, 31], fidelity susceptibility diverges near phase transitions or EPs [42, 43], and machine learning methods can be transferred from Hermitian to non-Hermitian systems [61]. In our work on entanglement Hamiltonians, we extend the Hermitian case to non-Hermitian systems and find that, for a nearly maximally entangled ground-state, the results remain consistent with the Hermitian case.
Acknowledgements
YCT is grateful to Chia-Yi Ju, Po-Yao Chang and Gunnar Möller for many invaluable discussions. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.
Funding information
YCT is grateful to the supports from National Science and Technology Council (NSTC) of Taiwan under grant No. 113-2112-M-A49-015-MY3.